the mean life of a sample of 400 fluorescent bulbs produced by a company is found to be 1570 hours with a standard deviation of 150hours. test the hypothesis that the mean life time of the bulbs produced by the company is 1600 hours at 1 percent level of significant.
1
Expert's answer
2019-10-21T13:05:18-0400

Given that "n=400, \bar{X}=1570, \sigma=150, \mu_0=1600"

The following null and alternative hypotheses need to be tested:

"H_0: \mu=1600"

"H_a:\mu\not=1600"

"n=400>30, \sigma" is known "=>z-test."

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a two-tailed test is "z_c=2.58"

The rejection region for this two-tailed test is "R=\lbrace z:|z|>2.58\rbrace."

The z-statistic is computed as follows:


"Z={\bar{X}-\mu_0 \over \sigma/\sqrt{n}}={1570-1600 \over 150/\sqrt{400}}=-4"

Since it is observed that "|z|=4>2.58=z_c", it is then concluded that the null hypothesis is rejected.  

Using the P-value approach:


"P(Z\leq-4)=0.00003167"

Two-tailed test

"2P(Z\leq -4)=0.00006334"

The p-value is p = 0.00006334 and since "p=0.00006334<0.01," it is concluded that the null hypothesis is rejected.

It is concluded that the null hypothesis "H_0" is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 1600, at the 0.01 significance level.

The 99% confidence interval is


"1550.681<\mu<1589.319"